Saccar Inc. expects to pay a dividend of $1.83 per share in one year. The current price of Saccar common stock is $30.04 per share. Flotation costs are $2.36 per share when the corporation issues new stock. What is the cost of internal common equity (retained earnings) if the long-term growth in dividends is projected to be 4 percent indefinitely? Submit your answer as a percentage and round to two decimal places (Ex. 0.00%)

Cantink Corporation plans to issue new common stock for a new expansion. The corporation’s existing common stock currently sells for $30.54. Management believes that they can issue new common stock at this price, incurring flotation costs of 2.3% of the current market price. What is the stock’s net market price (net proceeds)? Submit your answer as a dollar amount and round your answer to two decimal places (Ex. $0.00)
Answer #1


Solution: 01 rs=(D1/PO)+g rs=(1.83/30.04)+0.04 rs= 10.09% PO= market value per share rs = cost of internal equity g= growth r